The solubility of silver chloride AgCl is 1.2×10^-10 at 298 Kelvin. Calculate the solubility of silver chloride AgCl.

( Using ksp= s2)


1
Expert's answer
2022-03-09T12:11:03-0500

Solution:

AgCl(s) ⇌ Ag+(aq) + Cl(aq)

__S________S________S____

Thus,

S = S(AgCl) = [Ag+] = [Cl]


The Ksp expression for AgCl(s) is:

Ksp = [Ag+] × [Cl]

Therefore,

Ksp = [S] × [S]

1.2×10–10 = S2

S = (1.2×10–10)1/2 = 1.095×10–5

S(AgCl) = S = 1.095×10–5 M = 1.1×10–5 M

S(AgCl) = 1.1×10–5 M


Answer: The solubility of silver chloride (AgCl) is 1.1×10–5 M